OpenStudy (anonymous):

simplfy complex fraction: (X/4)-2 / (x/4)+2

OpenStudy (anonymous):

\[\frac{\frac{x}{4}-2}{\frac{x}{4}+2}\] multiply top and bottom by 4 to get \[\frac{x-8}{x+8}\]